6. Peggy-Sue loves to bake cookies and recently left her job at Cookie Monster, Inc., (making $125,000 per year) to start her own cookie-baking business. Her annual costs would include $40,000 for labor, $10,000 for rent, $35,000 for ingredients, and $5,000 for utilities. She withdrew $100,000 of her own savings to invest in the operation, which had been earning 10 percent interest per year. a. List Peggy-Sue’s total costs of operating her own cookie business, and specifically indicate which items are explicit costs and which are implicit costs. What are her total economic costs? (Yes, this part is a repeat from Chapter 11!) b. Peggy-Sue currently produces her own cookies and has revenues of $260,000 per year. Calculate both her accounting profit and economic profit. What do you think she should do?
Answer #1

a) TC = EXPLICIT+IMPLICIT

Explicit = labor+rent+ingredients+utilities = 40000+10000+35000+5000 = 90000

Implicit = foregone salary+10% of 100000 = 125000+10000 = 135000

Total cost = 90000+135000 = 225000

b) Accounting profit = total revenue - explicit costs = 260000-90000 = 170000

Economic profit = accounting profit - implicit costs = 170000-135000 = 35000

As Economic profit is positive, she should continue her business